# Real Noncommutative Convexity II: Extremality and nc convex functions

### v2

@@ Abstract.
We continue with the theory of real noncommutative (nc) convexity, following the recent and profound complex case developed by Davidson and Kennedy. The present paper focuses on the theory of nc extreme (and pure and maximal) points and the nc Choquet boundary in the real case, and on the theory of real nc convex and semicontinuous functions and real nc convex envelopes. Our main emphasis is on how these interact with complexification. For example some of our paper analyzes carefully how various notions of ‘extreme’ or ‘maximal’ interact with our earlier concept of the complexification of a convex set. Several new features appear in the real case, particularly in later sections of our paper, including the novel notion of the complexification of a nc convex function, and the complexification of the convex envelope of a nc function. With an Appendix by T. Russell.
@@ 1. Introduction
Very recently in [8] we developed the real case of the theory of real noncommutative (nc) convex sets in the sense of Davidson and Kennedy. We continue here with the theory of real noncommutative (nc) convexity, following the recent and profound complex case developed by those authors in the memoir [14]. As before, one of our primary goals is to develop the role and properties of the complexification in this context. Several motivations for our work are given in [8], one mentioned in more detail there is that ‘classical convexity’ is in some ways essentially a real theory, and so the real variant of the theory is likely to be important in the future in operator algebras, functional analysis, mathematical physics, and ‘quantum analysis’ (see e.g. [16] as a representative example of what part of this might look like). The present paper focuses on the real case of the topics from the middle sections of the deep memoir [14], namely on the theory of nc extreme (and pure and maximal) points and on Arveson’s nc Choquet boundary in the real case, and on the theory of real nc convex and semicontinuous functions and real nc convex envelopes; especially on how these interact with complexification. One use of this is as a tool to transfer results back and forth between the real and complex nc convexity theory, which often allows one to avoid difficulties. Some of our paper analyzes carefully how various notions of ‘extreme’ interact with our earlier concept of the complexification of a convex set. More sophisticated results, such as more advanced nc Choquet theory, will be discussed in a future paper. Several new features appear in the real case in later sections of our paper, including the novel notion of the complexification of a nc convex function on a nc convex set $K$, and the complexification of the convex envelope of a nc function on $K$. As just one example, we show that the convex envelope construction commutes with complexification. The convex envelope of a function $f$ defined on a convex set is of utmost importance in classical convexity theory, where it is sometimes called the ‘lower envelope’. (See e.g. [30]. Of course the upper/concave envelope may be treated via the convex envelope of f-f.) More generally we continue giving many foundational structural results for real nc convex sets and the functions on them.
@@
The real $C^{*}$-envelope of a real operator system, a.k.a. the noncommutative Shilov boundary, and most of its basic properties, were investigated some time ago (see e.g. the references at the start of [9, Section 5]), but in early sections of our paper we investigate its construction via the more recent ‘extremal’ and ‘nc Choquet boundary’ approach. We will see that the complexification of the nc Choquet boundary is incomparably more complicated than the complexification of the nc Shilov boundary, or of maximal elements or maximal ucp maps. That is, there are significant complications in the relationship between complexification and pure and extreme points of nc convex sets, boundary representations, etc. This is because the complexification of nc extreme points may not be nc extreme. (That this might be the case is not surprising in view of analogous results from the very recent paper [33], which works in the different setting of $C^{*}$-extreme maps on real $C^{*}$-algebras.) On the other hand maximal elements or maximal ucp maps behave perfectly in regard to complexification, and these are sufficient to yield the noncommutative Shilov boundary. From the outset it appeared that the real extremal theory might work out differently to the complex case, since it is so connected to irreducible representations, and irreps of real $C^{*}$-algebras work differently to the complex case in significant ways [29]. Fortunately however we find that there are no significant issues in this regard.
@@
There is of course a large and highly impressive literature (going back thirty years or so) on noncommutative or quantum versions of extreme points, Krein-Milman, etc, and also on noncommutative versions of convex functions. As just a few recent examples see for instance the remarkable papers [20, 22, 34, 35] (and the list of papers referred to on p. 2 of [8]), and references therein. (At the time of ArXiV submission of the present paper Scott McCullough also pointed out the very very recent [21].) Indeed there are even classical precursors e.g. in the operator (or matrix) convex and monotone functions going back to Löwner. Moreover this large existing theory of ‘operator/matrix convexity’ of which we speak often focuses on real examples (e.g. spectrahedra), and sometimes is even completely cast in the real framework, using ‘real’ tools. However we believe that there is no technical overlap of our work with this literature, since our goal here is mainly to extend the perspective of [14] to the real case, particularly emphasizing and developing the role of the complexification in this endeavor. (In some stretches of Section 2 where the complexification plays little part we are essentially just verifying and resolving potential complications in, and otherwise expositing, the real case of [14, Section 6]. Our contributions in this part of Section 2 are not intended to be novel, but rather to be useful and systematic for future ‘real users’.) Indeed aspects of the work earlier cited in this paragraph will hopefully interact very profitably with the matter of the present paper(s) in the future, and in any case this cited work certainly furnishes very many interesting and important real examples. Every complex nc convex set or complex nc convex function is a real one, but there are many interesting and important real examples which are not complex ones. Also, much of the work cited earlier in this paragraph is stated in the Wittstock-Effros-Webster-Winkler language [40] of matrix convexity. We refer readers to [14] for discussion of some of the differences between nc and matrix convexity.
@@
Turning to the structure of our paper, in Section 2 we begin by characterizing maximal elements in a complexification. This gives a new route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). We also discuss the real case of the theory of Davidson and Kennedy’s nc convexity variant of extreme, pure, and maximal points (which when ‘irreducible’ are essentially Arveson’s boundary representations). And of the associated nc Choquet and Shilov boundary, and its application to the nc Krein-Milman theorem and its converse. In most of the rest of this section we give a few points of clarification about the real case of the arguments, and note a few places where the arguments are different in the real case. In Section 3 we examine the connections between extreme, pure, and maximal points in the real and complex case, that is, how the these notions interact with the complexification of a nc convex set. One may view some of our work in this section as complementary to parts of the very recent paper [33]. In that paper some analogous results are established for $C^{*}$-extreme maps on real $C^{*}$-algebras. Every pure (and in particular any nc-extreme) ucp map on a $C^{*}$-algebra is $C^{*}$-extreme, but not vice versa. They do not treat general operator systems, thus their results would not apply to points in general nc convex sets. In Sections 4 and 5 we consider the real case of Davidson and Kennedy’s nc convex functions, including their multivalued nc convex functions, and convex envelopes. We give many foundational structural results here, for example developing the theory of the complexification in these settings and elucidating how it interacts with the basic objects and constructions. E.g. we show that the convex envelope construction commutes with complexification. As in [8], the complexification then allows us to derive the real variant of profound and extremely technical results from [14, Section 7] very quickly from the complex case.
@@
Finally, an Appendix by Travis Russell contains some details on the Dritschel-McCullough-Arveson maximal dilation, and on the converse to the Krein-Milman theorem, in the real case.
@@
As usual, many of the results for real operator systems here or in [8] apply or have variants for the larger class of real unital operator spaces: If $V$ is a real operator space then $V+V^{*},V\cap V^{*}$, and $\{x+x^{*}:x\in V\}$ are well defined real operator systems [9, 10].
@@
In keeping with the task and nature of our paper we will expect our readers to be a little familiar with, and reading alongside with parts of, [14] and our prequels [8, 9] (numbers in [14] refer to numbering in the v4 Arxiv version). Because of this we also do not need to be very pedantic or overly careful with definitions, preliminaries, or the history of the subject, which may usually be found referenced in more detail there. For general background on operator systems and spaces, and in particular on the definitions etc. in the rest of this section, we refer the reader to e.g. [36, 7, 14] and in the real case to e.g. [5, 9, 8]. It might also be helpful to also browse some of the other existing real operator space theory e.g. [38, 39, 10, 6]. Some basic real $C^{*}$- and von Neumann algebra theory may be found in [29].
@@
We write $M_{n}(\mathbb{R})$ for the real $n\times n$ matrices, or sometimes simply $M_{n}$ when the context is clear. Similarly in the complex case. It is important that $n$ is usually not necessarily an integer, but rather a cardinal bounded by a fixed cardinal $\kappa$. We sometimes use the quaternions $\mathbb{H}$ as an example: this is simultaneously a real operator system, a real Hilbert space, and a real C-algebra, usually thought of as a real *-subalgebra of $M_{4}(\mathbb{R})$ or $M_{2}(\mathbb{C})$. Its complexification is $M_{2}(\mathbb{C})$. The letters $H,L$ are usually reserved for real or complex Hilbert spaces. Every complex Hilbert space $H$ is a real Hilbert space, i.e. we forget the complex structure. More generally we write $X_{r}$ for a complex Banach space regarded as a real Banach space. We write $X_{\rm sa}$ for the selfadjoint elements in a *-vector space $X$. In the complex case $M_{n}(X)_{\rm sa}\cong(M_{n})_{\rm sa}\otimes X_{\rm sa}$, but this fails for real spaces. A subspace of $B(H)$ is unital if contains the identity, and a map $T$ is unital if $T(1)=1$. Our identities $1$ always have norm $1$. For $z\in M_{n}(\mathbb{C})$ we write $\text{Re }z$ for $x\in M_{n}(\mathbb{R})$ where $z=x+iy$ for $y\in M_{n}(\mathbb{R})$. Finally, for a cardinal $n$ we define the isometry $u_{n}=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$ where $1_{n}$ is the $n$-dimensional identity operator. We sometimes also write this as $u$.
@@
The natural morphisms between operator systems are unital completely positive (ucp) functions, which are linear maps $\varphi:V\to W$ that are unital and every amplification is positive (or equivalently selfadjoint and contractive). The isomorphisms (resp. embeddings) of operator systems which are used in this paper are bijective (resp. injective) ucp maps whose inverse (resp. inverse in its range) is ucp. These are called unital complete order isomorphisms (resp. unital complete order embeddings); or ucoi (resp. ucoe) for short.
Similarly a concrete operator space $E$ is a subspace of $B(H)$ with norms on $M_{n}(E)$ inherited from $B(H^{(n)})$. There are abstract characterizations of real/complex operator spaces and operator systems which we will not repeat here. If $E$ is a real operator system then $E$ can naturally be made into a complex operator system by complexification. The complexification $E_{c}$ of $E$ is the complex vector space $E_{c}$ consisting of elements $x+iy$ for $x,y\in E$. We give this a conjugate linear involution $(x+iy)^{*}=x^{*}-iy^{*}$. The matrix ordering $M_{n}(E_{c})^{+}$ will be defined by
@@
We also sometimes write $c(x+iy)$ for $c(x,y)$. We recall that this complexification is the unique one satisfying Ruan’s completely reasonable condition, namely that the map $\theta_{E}(x+iy)=x-iy$ is a ucoi. Note that for a real $C^{*}$-algebra $A$, and in particular for $A=M_{m}(\mathbb{R})$ the map $c:A_{c}\to M_{2}(A)$ is a faithful *-homomorphism.
In [8] we also defined the novel notion of the complexification of a nc convex set. A real nc convex set $K$ is a subset of $\mathcal{M}(E)$ for a real operator space $E$ that is graded, closed under direct sums, and closed under compressions. The set is compact if each level is compact in the point-weak topology. This is the same definition as in the complex case except that direct sums and compressions are done using real matrices. Equivalently, it is a set closed under nc convex combinations. The complexification of $K$, denoted by $K_{c}$, is the subset of $\mathcal{M}(E_{c})$ consisting of elements $x+iy$ such that $c(x,y)\in K$. Such expressions $c(x,y)$ or $c(x+iy)$ occur frequently in our paper(s). In [8] it is shown that $K_{c}$ is (the unique, under a very mild hypothesis) complex convex set containing $K$ such that $K$ generates $K_{c}$.
For a real operator space $E$, as in [14] we define $\mathcal{M}(E)=\bigsqcup_{n}\,M_{n}(E)$ (for $n$ cardinals bounded by some cardinal $\kappa$) with $M_{n}(E)$ the matrix space of $E$ for $X\subseteq\mathcal{M}(E)$. Here $\bigsqcup_{n}$ is the disjoint union. Define $X_{n}=X\bigcap M_{n}(E)$. In the case $E=\mathbb{R}$ write $\mathcal{M}=\mathcal{M}(\mathbb{R})$.
@@
We write $V_{r}$ for a complex operator system regarded as a real one. By a (real) (resp. complex) $C(X)$-space we will mean the continuous real (resp. complex) functions on a compact Hausdorff space $X$. Real commutative $C^{*}$-algebras are not all of this kind, for example $\mathbb{C}_{r}$ is not real *-isomorphic to a real $C(X)$-space. However real $C(X)$ spaces may be viewed as the real commutative $C^{*}$-algebras which have identity involution [24].

### v3

# arXiv:2507.22833v3[math.OA] 28 Feb 2026

@@ Abstract.
We continue the development of real noncommutative (nc) convexity, building on the recent and profound complex theory of Davidson and Kennedy. The present paper focuses on the theory of nc extreme points (and pure and maximal points) and the nc Choquet boundary in the real setting, as well as on the theory of real nc convex and semicontinuous functions and real nc convex envelopes. Our main emphasis is on how these notions interact with complexification. In particular, parts of the paper analyze in detail how various notions of “extreme” or “maximal” relate to our earlier concept of the complexification of a convex set. Several new features emerge in the real case, especially in the later sections, including the novel notions of the complexification of a nc convex function and of the complexification of the convex envelope of a nc function. With an appendix by T. Russell.
@@ 1. Introduction
Very recently, in [8], we developed the real case of the theory of noncommutative (nc) convex sets in the sense of Davidson and Kennedy. In the present work, we continue the development of real nc convexity, guided by the recent and profound complex theory developed by those authors in the memoir [16]. As before, one of our primary goals is to understand the role and properties of complexification in this setting. Several motivations for our work are described in [8]. One, discussed there in more detail, is that classical convexity is, in many respects, essentially a real theory. Consequently, the real variant of nc convexity is likely to play an important role in future developments in operator algebras, functional analysis, mathematical physics, and quantum analysis (see for example [18] as a representative example of what part of this might look like). The present paper focuses on the real analogues of topics corresponding to the middle sections of the memoir [16]. This includes the theory of nc extreme points (and pure and maximal points), Arveson’s nc Choquet boundary in the real case, and the theory of real nc convex and semicontinuous functions and real nc convex envelopes. In particular we emphasize how all of these interact with complexification. One use of this is as a tool for transferring results between the real and complex theories of nc convexity, often allowing one to circumvent technical difficulties. Portions of the paper analyze in detail how various notions of “extreme” interact with our earlier notion of the complexification of a convex set. More sophisticated developments, such as deeper aspects of nc Choquet theory, will be addressed in future work.
Several new features arise in the real case, particularly in the later sections of the paper. These include the novel notions of the complexification of a nc convex function on a nc convex set $K$, and of the complexification of the convex envelope of a nc function on $K$. As just one example, we show that the convex envelope construction commutes with complexification. The convex envelope of a function $f$ defined on a convex set is of central importance in classical convexity theory, where it is sometimes referred to as the “lower envelope” (see, for example, [33]; the concave or upper envelope may of course be treated via the convex envelope of f-f. More generally, we continue to develop foundational structural results for real nc convex sets and the nc functions defined on them.
@@
The real $C^{*}$-envelope of a real operator system, a.k.a. the noncommutative Shilov boundary, and most of its basic properties, were investigated some time ago (see e.g. the references at the start of [9, Section 5]). In early sections of our paper we investigate its construction via the more recent ‘extremal’ and ‘nc Choquet boundary’ approach. We will see that the complexification of the nc Choquet boundary is incomparably more complicated than the complexification of the nc Shilov boundary, or of maximal elements or maximal ucp maps. That is, there are significant complications in the relationship between complexification and pure or extreme points of nc convex sets, boundary representations, and related notions. This arises because the complexification of a nc extreme point need not remain nc extreme. That such behavior can occur is not surprising in light of analogous results in the recent paper [36], which treats the different setting of $C^{*}$-extreme maps on real $C^{*}$-algebras. By contrast, maximal elements—or maximal unital completely positive maps—behave perfectly under complexification, and these are sufficient to recover the noncommutative Shilov boundary. From the outset, it appeared plausible that the real extremal theory might differ substantially from the complex case, given its close connection to irreducible representations and the fact that irreducible representations of real $C^{*}$-algebras differ in important ways from their complex counterparts [32]. Fortunately, we find that no significant difficulties arise in this regard.
@@
There is of course a large and highly impressive literature (going back thirty years or so) on noncommutative or quantum versions of extreme points, Krein-Milman, etc, and also on noncommutative versions of convex functions. As just a few recent examples see for instance the remarkable papers [22, 24, 37, 38] (and the list of papers referred to on p. 2 of [8]), and references therein. (At the time of ArXiV submission of the present paper Scott McCullough also pointed out the very recent [23].) Indeed there are even classical precursors, e.g. in the operator (or matrix) convex and monotone functions going back to Löwner. Moreover this large existing theory of ‘operator/matrix convexity’ of which we speak, often focuses on real examples (e.g. spectrahedra), and sometimes is even completely cast in the real framework, using ‘real’ tools. However we believe that there is no technical overlap of our work with this literature, since our goal here is mainly to extend the perspective of [16] to the real case, particularly emphasizing and developing the role of the complexification in this endeavor. (In some stretches of Section 2 where the complexification plays little part we are essentially just verifying and resolving potential complications in, and otherwise expositing, the real case of [14, Section 6]. Our contributions in this part of Section 2 are not intended to be novel, but rather to be systematic and useful for future ‘real users’.) Indeed aspects of the work earlier cited in this paragraph will hopefully interact in the future very profitably with the matter of the present paper(s), and in any case this cited work certainly furnishes very many interesting and important real examples. Every complex nc convex set or complex nc convex function is a real one, but there are many interesting and important real examples which are not complex ones. Also, much of the work cited earlier in this paragraph is stated in the Wittstock-Effros-Webster-Winkler language [43] of matrix convexity. We refer readers to [16] for discussion of some of the differences between nc and matrix convexity.
@@
Turning to the structure of our paper, in Section 2 we begin by characterizing maximal elements in a complexification. This gives a new route to the construction of the $C^{*}$-envelope of a real operator system (or unital operator space). We also discuss the real case of the theory of Davidson and Kennedy’s nc convexity variant of extreme, pure, and maximal points (which where ‘irreducible’ are essentially Arveson’s boundary representations). And of the associated nc Choquet and Shilov boundary, and its application to the nc Krein-Milman theorem and its converse. In most of the rest of Section 2 we give a few points of clarification about the real case of the arguments, and note a few places where the arguments are different in the real case. In Section 3 we examine the connections between nc extreme, pure, and maximal points in the real and complex case, that is, how the these notions interact with the complexification of a nc convex set. One may view some of our work in this section as complementary to parts of the very recent paper [36]. In that paper some analogous results are established for $C^{*}$-extreme maps on real $C^{*}$-algebras. Every pure (and in particular any nc-extreme) ucp map on a $C^{*}$-algebra is $C^{*}$-extreme, but not vice versa. They do not treat general operator systems, thus their results would not apply to points in general nc convex sets.
In Sections 4 and 5 we consider the real case of Davidson and Kennedy’s nc convex functions, including their multivalued nc convex functions, and convex envelopes. We give many foundational structural results here, for example developing the theory of the complexification in these settings and elucidating how it interacts with the basic objects and constructions. E.g. we show that the convex envelope construction commutes with complexification. As in [8], the complexification then allows us to derive the real variant of profound and extremely technical results from [16, Section 7] very quickly from the complex case.
@@
Finally, an Appendix by Travis Russell contains some details on the converse to the Krein-Milman theorem, in the real case.
@@
As usual, many of the results for real operator systems here or in [8] apply or have variants for the larger class of real unital operator spaces: If $V$ is a real unital operator space then $V+V^{*},V\cap V^{*}$, and $\{x+x^{*}:x\in V\}$ are well defined real operator systems [9, 10].
@@
In keeping with the task and nature of our paper we will expect our readers to be a little familiar with, and reading alongside with parts of, [16] and our prequels [8, 9] (numbers in [16] refer to numbering in the v4 Arxiv version). Because of this we do not need to be very pedantic or overly careful with definitions, preliminaries, or the history of the subject, which may usually be found referenced in more detail there. For general background on operator systems and spaces, and in particular on the definitions etc. in the rest of this section, we refer the reader to e.g. [39, 7, 16], and in the real case to e.g. [5, 9, 8]. It might also be helpful to also browse some of the other existing real operator space theory e.g. [41, 42, 10, 6]. Some basic real $C^{*}$- and von Neumann algebra theory may be found in [32].
@@
We write $M_{n}(\mathbb{R})$ for the real $n\times n$ matrices, or sometimes simply $M_{n}$ when the context is clear. Similarly in the complex case. We write ${\mathfrak{R}}\,a$ for $\frac{1}{2}(a+a^{*})$, while for $z\in M_{n}(\mathbb{C})$ we write $\text{Re}\,z$ for $x\in M_{n}(\mathbb{R})$ where $z=x+iy$ for $y\in M_{n}(\mathbb{R})$. We sometimes use the quaternions $\mathbb{H}$ as an example: this is simultaneously a real operator system, a real Hilbert space, and a real C-algebra, usually thought of as a real *-subalgebra of $M_{4}(\mathbb{R})$ or $M_{2}(\mathbb{C})$. Its complexification is $M_{2}(\mathbb{C})$. The letters $H,L$ are usually reserved for real or complex Hilbert spaces. Every complex Hilbert space $H$ is a real Hilbert space, i.e. we may forget the complex structure. More generally we write $X_{r}$ for a complex Banach space regarded as a real Banach space. We write $X_{\rm sa}$ for the selfadjoint elements in a *-vector space $X$. In the complex case $M_{n}(X)_{\rm sa}\cong(M_{n})_{\rm sa}\otimes X_{\rm sa}$, but this fails for real spaces. A subspace of $B(H)$ is unital if contains the identity, and a map $T$ is unital if $T(1)=1$. Our identities $1$ always have norm $1$. For $z\in M_{n}(\mathbb{C})$ we write $\text{Re}\,z$ for $x\in M_{n}(\mathbb{R})$ where $z=x+iy$ for $y\in M_{n}(\mathbb{R})$. For a cardinal $n$ we define the isometry $u_{n}=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$ where $1_{n}$ is the $n$-dimensional identity operator. We sometimes often write this as $u$ (suppressing the subscript).
@@
The natural morphisms between operator systems are unital completely positive (ucp) functions, which are linear maps $\varphi:V\to W$ that are unital and every amplification is positive (or equivalently selfadjoint and contractive). The isomorphisms (resp. embeddings) of operator systems which are used in this paper are bijective (resp. injective) ucp maps whose inverse (resp. inverse in its range) is ucp. These are called unital complete order isomorphisms (resp. unital complete order embeddings).
Similarly, a concrete operator space $E$ is a subspace of $B(H)$ with norms on $M_{n}(E)$ inherited from $B(H^{(n)})$. There are abstract characterizations of real/complex operator spaces and operator systems which we will not repeat here. If $E$ is a real operator system then $E$ can naturally be made into a complex operator system by complexification. The complexification $E_{c}$ of $E$ is the complex vector space $E_{c}$ consisting of elements $x+iy$ for $x,y\in E$. We give this a conjugate linear involution $(x+iy)^{*}=x^{*}-iy^{*}$. The matrix ordering $M_{n}(E_{c})^{+}$ will be defined by
@@
We also sometimes write $c(x+iy)$ for $c(x,y)$. We recall that this complexification is the unique one satisfying Ruan’s completely reasonable condition, namely that the map $\theta_{E}(x+iy)=x-iy$ is a unital complete order isomorphism. Note that for a real $C^{*}$-algebra $A$, and in particular for $A=M_{m}(\mathbb{R})$ the map $c:A_{c}\to M_{2}(A)$ is a faithful *-homomorphism.
For a real or complex operator space $E$ and a possibly infinite cardinal, $M_{n}(E)$ is the space of matrices whose ‘finitely supported’ submatrices have uniformly bounded norm. In the case $E=\mathbb{F}$ write $M_{n}=M_{n}(\mathbb{F})$. Thus $M_{n}$ is $B(\ell^{2}_{n})$. Indeed for every Hilbert space $H$, $B(H)\cong M_{n}$ *-isomorphically for some $n$, after one chooses an orthonormal basis. For a real operator space $E$, as in [16] we define $\mathcal{M}(E)=\bigsqcup_{n}\,M_{n}(E)$ (for $n$ cardinals bounded by some large enough cardinal $\kappa$) with $M_{n}(E)$ the matrix space of $E$ for $X\subseteq\mathcal{M}(E)$. Here $\bigsqcup_{n}$ is the disjoint union. Define $X_{n}=X\bigcap M_{n}(E)$. In the case $E=\mathbb{R}$ write $\mathcal{M}=\mathcal{M}(\mathbb{R})$. In [8] we also defined the novel notion of the complexification of a nc convex set. (This has already found applications, for example in [11].) A real nc convex set $K$ is a subset of $\mathcal{M}(E)$ for a real operator space $E$ that is graded, closed under direct sums, and closed under compressions. The set is compact if each level is compact in the point-weak topology. This is the same definition as in the complex case except that direct sums and compressions are done using real matrices. Equivalently, it is a set closed under nc convex combinations. The complexification of $K$, denoted by $K_{c}$, is the subset of $\mathcal{M}(E_{c})$ consisting of elements $x+iy$ such that $c(x,y)\in K$. Such expressions $c(x,y)$ or $c(x+iy)$ occur frequently in our paper(s). Similarly, for a cardinal $n$ we define as in [8] the isometry $u_{n}=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$ where $1_{n}$ is the $n$-dimensional identity operator. We sometimes also write this as $u$. In [8] it is shown that $K_{c}$ is (the unique, under a very mild hypothesis) complex convex set containing $K$ such that $K$ generates $K_{c}$.
@@
We write $V_{r}$ for a complex operator system regarded as a real one. By a (real) (resp. complex) $C(X)$-space we will mean the continuous real (resp. complex) functions on a compact Hausdorff space $X$. Real commutative $C^{*}$-algebras are not all of this kind, for example $\mathbb{C}_{r}$ is not real *-isomorphic to a real $C(X)$-space. Real $C(X)$ spaces may be viewed as the real commutative $C^{*}$-algebras which have identity (i.e. trivial) involution [26].